OpenStudy (anonymous):

Write this mixed number as a percent. 1 27/50

OpenStudy (unklerhaukus):

\[1~{ 27 \over50}\]\[={50\over 50} +{27 \over 50}\]\[=77/50\]\[=154/100\]\[=154\%\]